3.Find the area of the region bounded by the parametric curve and the x-axis. (10 pts) = 6 (0- sin 0) y=6(1 - cos 0) 0<02T
Find the slope of the tangent line at the given point. (10 pts) 4. r 2+sin 30, 0=T/4
Answer #1

- Area bounded by Parametric wave & xani n = 6(0-sino) OCOLRI da= 611-coso) do y = 6(1-coso) = Tych = 6-6 coso) 6(1-coso) do